From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon May 05 2003 - 23:55:14 EDT
------------------------------------------------------------ Date: Mon May 5 23:55:13 2003, Shift: 16:00-24:00 Supervisor: kh, with: Log book#14 pages 96 - 96 Shift Summary: ------------------------------------------------------------ Run 8827 in progress at shift start; Store ended at 18:00 Another store started at 20:00. Took data while they were steering. Got 100k before they said they would dump the beam without even releasing it. Runs accomplished Run trig 3 trig 6 8827 99K 46K FS 1/8B Change FS magnet polarity to A 8828 69K 7.2K FS 1/2A 8834 100K 46K FS 1/8A No detector problems detected. ------------------------------------------------------------
This archive was generated by hypermail 2.1.5 : Mon May 05 2003 - 23:59:01 EDT